OpenStudy (anonymous):

The digits 1,2,3,4,5, and 6 are each used once to compose a six digit number abcdef, such that the three digits number abc is divisible by 5, cbe is divisible by 3, and def is divisible by 11. The digit a is?

OpenStudy (anonymous):

!!

OpenStudy (anonymous):

a = 2.

OpenStudy (anonymous):

Does e = 4 b=6 c=5 d=3 f=1

OpenStudy (anonymous):

Satisfy, right?

OpenStudy (anonymous):

how did you get that?

OpenStudy (mathmagician):

the numbers, that satisfies these conditions, are: 645132 265143 645231 315264 265341 315462

OpenStudy (anonymous):

Okay. Condition one : abc divisible by 5. Implies c = 5 Now 5 + b + e must be a multiple of 3. Now make cases. Which is quite easy here. Max multiple of three possible = 6 + 5+4 = 15 Therefore cases are. 5 b e 4 6 / 6 4 ( When total is 15, b + e = 10) 6 1 / 1 6 ( total = 12 , b + e =7) 5 2 / 2 5 ( ") 4 3 / 3 4 (") ---- Less than that - Not three digit.

OpenStudy (anonymous):

Also e = 5 -- Not possible ( c is 5 - One case eliminated) Now condition 3, d + f - e = Multiple of 11. Has to be 0, As 11-- Not possible. Therefore d + f = e. Possible values of e = 6,4,1,2, 1 , 2 --- Not possible as sum of two numbers from given list.

OpenStudy (anonymous):

And 3^

OpenStudy (anonymous):

Yes. That gives all the cases mathmagician drew up :)

OpenStudy (anonymous):

thank!